T(n)=8+7(n-1). N=0
Answer by 119078(26)

Now something to remember is n is like x and t(n) is like f(x) or y so it is easier to do that way change it. I will have it using x and y at the bottom.
Start by plugging in 0 for n
T(0)=8+7(0-1)
distribute 7 into (0-1) which gives you +0-7. Because it is a zero you can just drop it and no one be the wiser! :D
T(0)=8-7
Next add 8 and -7 and then you have what T(n) is when n is 0.
T(0)=1
Now for with x and y the numbers and way to do it is the same just using different variables.
y=8+7(x-1) x=0
y=8+7(0-1)
y=8-7
y=1
I hope seeing it both ways helped you out.